# Can a calibration metric be both testable and actionable?

### v1

@@ 4.1 Testability of dCE and Cutoff Calibration Error
We will next see that dCE and Cutoff Calibration Error are both testable—there are consistent estimators for $\Delta_{\operatorname{dCE}}(f)$ or $\Delta_{\operatorname{Cutoff}}(f)$, and consequently it is possible to test the hypothesis $\Delta_{\operatorname{dCE}}(f)\leq c$ or $\Delta_{\operatorname{Cutoff}}(f)\leq c$. In particular, as we will formalize below, we can construct practical procedures to return a function $f_{n}$ guaranteed to satisfy the metric.
First we consider Cutoff Calibration Error.
@@
$\widehat{\Delta}_{\operatorname{Cutoff}}(f)=\sup_{I:\text{ interval}}\left|\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-f(X_{i})){{\mathbf{1}_{\left\{f(X_{i})\in I\right\}}}}\right|.$ (13)
@@
$\left|\widehat{\Delta}_{\operatorname{Cutoff}}(f)-\Delta_{\operatorname{Cutoff}}^{P}(f)\right|\leq\frac{20+\sqrt{2\log(1/\delta)}}{\sqrt{n}}$ (14)
Now let us formalize how this leads to a practical procedure. Given a data set $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})\overset{iid}{\sim}P$, a threshold $c>0$ and $\delta\in(0,1)$, and a model fitting algorithm $\mathcal{A}$ (e.g., a neural net),
@@
In particular, note that the result holds uniformly over any distribution $P$, and the procedure does not require any knowledge of $P$—that is, this is a distribution-free guarantee. Moreover, there are no restrictions on the model fitting algorithm $\mathcal{A}$, and so the returned model $f_{n}$ may be quite complex and adaptive to the data, if we choose our algorithm well for the task at hand.
Next, consider dCE. In fact, essentially the same result holds: Błasiok et al. (2023, Section 9) establishes that dCE can also be estimated using an i.i.d. sample, with error $\mathcal{O}_{P}(n^{-1/2})$, and therefore, the same style of two-stage procedure would lead to an analogous guarantee.
@@ 4.2 Hardness result: ECE is not testable
We can think of this quantity as an “effective variability” of $P$—in particular if $\textnormal{Var}(Y\mid X)\geq\sigma^{2}>0$ almost surely, then $\sigma^{2}_{\gamma}(P)\geq\sigma^{2}$ for any $\gamma$. Next, we also define a notion of “effective support size”:
@@
or $S_{\gamma}(f,P)=+\infty$ if the set is empty. If $f(X)$ takes at most $k$ many values (under $X\sim P$), then $S_{\gamma}(f,P)\leq k$ for any $\gamma$ [^3] As for calibration error, when the function $f_{n}$ is data-dependent, we treat $f_{n}$ as fixed for the purpose of calculating $S_{\gamma}(f_{n},P)$—that is, $\mathbb{P}_{P}(f_{n}(X)\in\{t_{1},\dots,t_{k}\})$ should be interpreted as a probability calculated with respect to an independent draw of $X$, when we condition on the trained model $f_{n}$.
@@
$\mathcal{A}:(\mathcal{X}\times[0,1])^{n}\rightarrow\{\text{ measurable functions }f:\mathcal{X}\rightarrow[0,1]\}$ (15)
@@
If $c$ and $\delta$ are small (i.e., $\mathcal{A}$ offers a guarantee on ECE), then for any $P$ with an effective variance bounded away from zero, we must have $S_{\gamma}(f_{n},P)\leq n^{2}$ with high probability—that is, $f_{n}$ must be a function that is piecewise constant on most of the domain, with effective support at most $\mathcal{O}(n^{2})$.
@@ 5.1 Isotonic regression enjoys uniform asymptotic control of Cutoff Calibration Error
Suppose $(X_{i},Y_{i})\overset{iid}{\sim}$ for $i=1,\dots,n$. Assume $Y_{i}\in[0,1]$ and $f$ is fixed. Let $h_{n}$ refer to running isotonic regression on $\{(f(X_{i}),Y_{i})\}_{i=1}^{n}$. Fix $\delta>0$. Then,
$\Delta_{\operatorname{Cutoff}}^{P}(h_{n}\circ f)\leq\frac{30+2\sqrt{2\log(2/\delta)}}{\sqrt{n}}$ (16)
@@ 5.2 Platt scaling does not necessarily calibrate, even in dCE
Platt scaling (Platt et al., 1999) is a popular post-hoc calibration method which essentially consists of running a univariate logistic regression of observed $Y$ values on previous forecasts $f(X)$. [^4] For finite samples, Platt scaling minimizes logistic loss with $\tilde{Y}_{i}=Y_{i}\cdot\frac{\sum_{i=1}^{n}Y_{i}+1}{\sum_{i=1}^{n}Y_{i}+2}+(1-Y_{i})\frac{1}{\sum_{i=1}^{n}(1-Y_{i})+2}$ in place of $Y_{i}$ to improve empirical performance. The issue, in short, is that there is no guarantee that Platt scaling will converge asymptotically to a calibrated function. Since $\Delta_{\operatorname{dCE}}(f)=0$ if and only $f$ is perfectly calibrated, this implies that Platt scaling cannot provide strong asymptotic calibration on dCE, ECE, and Cutoff Calibration Error. In Figure 4, we exhibit an example of a data-generation process for which the population Platt scaling algorithm yields an $h:[0,1]\rightarrow[0,1]$ (attained by optimizing logistic loss in expectation instead of over samples) that has non-zero dCE: $\Delta_{\operatorname{dCE}}(h\circ f)>0$.
@@ 6 Discussion
There are several promising research directions and intriguing connections that stem from our findings. First, while our work focuses on the i.i.d. setting, we think our findings may hold ramifications for the branch of calibration research that focuses on sequential prediction and could potentially inform the design of new online calibration algorithms. Second, we believe our research bears connections to research on omniprediction (Gopalan et al., 2021), which shows that multicalibration (Hébert-Johnson et al., 2018), a concept closely related to weighted calibration error, ensures simultaneous optimization across the class of loss functions that are convex and Lipschitz. Third, there may be connections between our work and Sahoo et al. (2021), who also focus on “threshold” decision rules, although in a different setting than us, where forecasters provide a CDF estimate for a real-valued output. Fourth, the monotonic constraint used in our decision-theoretic guarantees may provide intriguing results in other contexts, such as that of Roth and Shi (2024), who examine the impact of the dimensionality of a prediction space on swap regret and calibration. Finally, while the decision-theoretic guarantees of Cutoff Calibration Error are broadest when $Y\in\{0,1\}$, we believe that there may be a more general guarantee related to Cutoff Calibration Error when $Y\in[0,1]$.
@@ 7 Acknowledgments
The authors gratefully acknowledge the National Science Foundation via grant DMS-2023109. J.S. and R.F.B. were partially supported by the Office of Naval Research via grant N00014-24-1-2544. J.S. was also partially supported by the Margot and Tom Pritzker Foundation. Z.R. was supported by NSF grant DMS-2413135. R.M.W. was partially supported by the NSF-Simons National Institute for Theory and Mathematics in Biology (NITMB) via grants NSF (DMS-2235451) and Simons Foundation (MP-TMPS-00005320).
@@ Appendix A Relationships among calibration measures: proofs and examples
$\mathbb{E}|f(X)-g(X)|\leq\Delta_{\operatorname{dCE}}(f)+\epsilon.$
@@
Next, let $w$ be any function taking values in $[-1,1]$. We need to bound $\mathbb{E}[w(f(X))\cdot(Y-f(X))]$. First, define
@@
$\mathbb{P}\{w(f(X))\neq\tilde{w}(g(X))\}\leq\frac{1}{b_{f}}\cdot\mathbb{E}{|f(X)-g(X)|}\leq\frac{1}{b_{f}}\left(\Delta_{\operatorname{dCE}}(f)+\epsilon\right).$
Finally, we prove ECE. We have
@@
$\Delta_{\operatorname{Cutoff}}(f)\leq\Delta_{\operatorname{ECE}}(f)$ (17)
@@
    This follows immediately from expressing ECE as an instantiation of wCE. ∎
@@
$\Delta_{\operatorname{dCE}}(f)^{2}/9\leq\Delta_{\operatorname{Cutoff}}(f)$ (18)
@@
    $\psi(z):=\sum_{j=1}^{N}j{{\mathbf{1}_{\left\{z\in A_{j}\right\}}}}$ (20)
@@
    $\mathbb{E}[Y\mid g(X)]=\mathbb{E}[\mathbb{E}[Y\mid\psi(f(X))]\mid g(X)]=\mathbb{E}[g(X)\mid g(X)]=g(X)$ (22)
@@
    $\psi(f(X))=j\implies f(X),V\in A_{j}\implies|f(X)-V|\leq\mathrm{diam}(A_{j})=1/N$ (24)
@@
$\mathbb{E}|g(X)-f(X)|\leq\mathbb{E}|g(X)-V|+\mathbb{E}|V-f(X)|\leq\mathbb{E}|g(X)-V|+1/N$ (25)
@@
$\mathbb{E}|g(X)-V|$ $=\mathbb{E}|\mathbb{E}[Y-f(X)\mid\psi(f(X))]|$ (26)
$=\sum_{j=1}^{N}\mathbb{P}\{f(X)\in A_{j}\}|\mathbb{E}[(Y-f(X))\mid f(X)\in A_{j}]|$ (27)
$=\sum_{j=1}^{N}|\mathbb{E}[(Y-f(X)){{\mathbf{1}_{\left\{f(X)\in A_{j}\right\}}}}]|$ (28)
$\leq N\Delta_{\operatorname{Cutoff}}(f)$ (29)
@@
$\Delta_{\operatorname{dCE}}(f)\leq N\Delta_{\operatorname{Cutoff}}(f)+1/N$ (30)
@@
$\Delta_{\operatorname{dCE}}(f)\leq 2\sqrt{\Delta_{\operatorname{Cutoff}}(f)}+\Delta_{\operatorname{Cutoff}}(f)\leq 3\sqrt{\Delta_{\operatorname{Cutoff}}(f)}$ (31)
@@
    The first inequality follows immediately from observing that indicator functions over intervals have a total variation of at most 2 and that, if $g\in\mathcal{B}_{M}$, then $-g\in\mathcal{B}_{M}$
    For the second inequality, by Lemma A.3, $\exists g_{1},g_{2}:[0,1]\rightarrow[0,\frac{M+1}{2}]$ non-decreasing such that $g=g_{1}-g_{2}$.
[…]
@@
$\sup_{g\in\mathcal{B}_{M}}\mathbb{E}[(Y-f(X))g(f(X))]$ (32)
$\leq\sup_{g_{1},g_{2}\in\mathcal{G}}\left(\mathbb{E}[(Y-f(X))g_{1}(f(X))]-\mathbb{E}[(Y-f(X))g_{2}(f(X))]\right)$ (33)
$=\sup_{g\in\mathcal{G}}\mathbb{E}[(Y-f(X))g(f(X))]+\sup_{g\in\mathcal{G}}\mathbb{E}[(Y-f(X))(-g(f(X)))]$ (34)
$\leq 2\sup_{g\in\mathcal{G}\cup(-\mathcal{G})}\mathbb{E}[(Y-f(X))g(f(X))]$ (35)
$=2\cdot\sup_{g\in\mathcal{G}^{\prime}\cup(-\mathcal{G}^{\prime})}\mathbb{E}[(Y-f(X))((M+1)\cdot g(f(X))-0.5(M+1))]$ (36)
$\leq 3\cdot(M+1)\sup_{g\in\mathcal{G}^{\prime}\cup(-\mathcal{G}^{\prime})}\mathbb{E}[(Y-f(X))g(f(X))]$ (37)
$\leq 3\cdot(M+1)\Delta_{\operatorname{Cutoff}}(f)$ (38)
@@
$\mathcal{G}^{\prime}=\overline{\mathrm{conv}}\{g(x)={{\mathbf{1}_{\left\{x\in[t,1]\right\}}}}\mid t\in[0,1]\}$ (39)
@@
$\sup_{g\in\mathcal{G}^{\prime}\cup(-\mathcal{G}^{\prime})}\mathbb{E}[(Y-f(X))g(f(X))]$ (40)
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ =\max\left\{\sup_{g\in\mathcal{G}^{\prime}}\mathbb{E}[(Y-f(X))g(f(X))],\sup_{g\in\mathcal{G}^{\prime}}\mathbb{E}[(Y-f(X))(-g(f(X)))]\right\}$ (41)
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ =\max\left\{\sup_{t}\mathbb{E}[(Y-f(X)){{\mathbf{1}_{\left\{f(X)\in[t,1]\right\}}}}],\sup_{t}\mathbb{E}[(Y-f(X))(-{{\mathbf{1}_{\left\{f(X)\in[t,1]\right\}}}})]\right\}$ (42)
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leq\sup_{t}|\mathbb{E}[(Y-f(X)){{\mathbf{1}_{\left\{f(X)\in[t,1]\right\}}}}]|$ (43)
@@
Suppose $f:[0,1]\rightarrow[0,1]$ and has total variation at most $M$. Then, $\exists f_{1},f_{2}:[0,1]\rightarrow[0,\frac{M+1}{2}]$ non-decreasing such that $f=f_{1}-f_{2}$
@@
    To show $f_{1}$ is non-decreasing, let $x\leq y$, where $x,y\in[0,1]$

### v2

@@ 4.1 Testability of dCE and Cutoff Calibration Error
We will next see that Cutoff Calibration Error and dCE are both testable: it is possible to test the hypothesis $\Delta_{\operatorname{Cutoff}}(f)\leq c$ or $\Delta_{\operatorname{dCE}}(f)\leq c$. In particular, as we will formalize below, we can construct practical procedures to return a function $f_{n}$ guaranteed to satisfy a bound on calibration error.
We begin by considering Cutoff Calibration Error. First, we establish that the natural estimator for $\Delta_{\operatorname{Cutoff}}(f)$ is consistent.
@@
$\widehat{\Delta}_{\operatorname{Cutoff}}(f)=\sup_{I:\text{ interval}}\left|\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-f(X_{i})){{\mathbf{1}_{\left\{f(X_{i})\in I\right\}}}}\right|.$
@@
$\left|\widehat{\Delta}_{\operatorname{Cutoff}}(f)-\Delta_{\operatorname{Cutoff}}(f)\right|\leq\frac{20+\sqrt{2\log(1/\delta)}}{\sqrt{n}}.$
In particular, we can therefore use this estimator to test whether cutoff calibration error is low. In other words, given some threshold $c>0$, if $\widehat{\Delta}_{\operatorname{Cutoff}}(f)\leq c-\frac{20+\sqrt{2\log(1/\delta)}}{\sqrt{n}}$ then we can certify (with $1-\delta$ confidence) that $\Delta_{\operatorname{Cutoff}}(f)\leq c$—and in particular, this procedure may have nontrivial power as soon as $n\gtrsim\frac{\log(1/\delta)}{c^{2}}$. An analogous result on testability for a closely related calibration measure is provided by Okoroafor et al. (2025, Lemma C.2).
More generally, however, we may be interested in providing a procedure that is guaranteed to offer Cutoff Calibration Error bounded by some threshold $c$, rather than testing whether Cutoff Calibration Error is bounded by $c$ (or not). We can also use the result above to provide a procedure that offers such a guarantee. Given a data set $(X_{1},Y_{1}),\dots,(X_{n},Y_{n})\overset{iid}{\sim}P$, a threshold $c>0$ and $\delta\in(0,1)$, and a model fitting algorithm $\mathcal{A}$ (e.g., a neural net),
@@
In particular, note that the result holds uniformly over any distribution $P$, and the procedure does not require any knowledge of PP—that is, this is a distribution-free guarantee. Moreover, there are no restrictions on the model fitting algorithm $\mathcal{A}$, and so the returned model $f_{n}$ may be quite complex and adaptive to the data, if we choose our algorithm well for the task at hand.
Next, consider dCE. In fact, since $\Delta_{\operatorname{dCE}}(f)\leq 3\sqrt{\Delta_{\operatorname{Cutoff}}(f)}$ by Proposition 2.1, this means that any guarantee of an upper bound on $\Delta_{\operatorname{Cutoff}}(f)$ (as in Corollary 4.1 above, say) immediately yields a bound on $\Delta_{\operatorname{dCE}}(f)$, as well. Relatedly, Błasiok et al. (2023, Section 9) establish that a calibration measure related to dCE can be estimated consistently using an i.i.d. sample, with error $\mathcal{O}_{P}(n^{-1/2})$, which leads to lower and upper bounds on dCE, as well.
@@ 4.2 Hardness result: ECE is not testable
We can think of this quantity as an “effective variability” of PP—in particular if $\textnormal{Var}(Y\mid X)\geq\sigma^{2}>0$ almost surely, then $\sigma^{2}_{\gamma}(P)\geq\sigma^{2}$ for any $\gamma$. Next, we also define a notion of “effective support size”:
@@
or $S_{\gamma}(f,P)=+\infty$ if the set is empty. If $f(X)$ takes at most $k$ many values (under $X\sim P$), then $S_{\gamma}(f,P)\leq k$ for any $\gamma$. [^5] As for calibration error, when the function $f_{n}$ is data-dependent, we treat $f_{n}$ as fixed for the purpose of calculating $S_{\gamma}(f_{n},P)$—that is, $\mathbb{P}_{P}(f_{n}(X)\in\{t_{1},\dots,t_{k}\})$ should be interpreted as a probability calculated with respect to an independent draw of $X$, when we condition on the trained model $f_{n}$.
@@
$\mathcal{A}:(\mathcal{X}\times[0,1])^{n}\rightarrow\{\text{ measurable functions }f:\mathcal{X}\rightarrow[0,1]\}$
@@
If $c$ and $\delta$ are small (i.e., $\mathcal{A}$ offers a meaningful guarantee on ECE), then for any $P$ with an effective variance bounded away from zero, we must have $S_{\gamma}(f_{n},P)\leq n^{2}$ with high probability—that is, $f_{n}$ must be a function that is piecewise constant on most of the domain, with effective support at most $\mathcal{O}(n^{2})$.
@@
Estimating ECE via binning?
In practice, it is common to estimate ECE with a binning-based approach: that is, given a function $f$ and a partition $0,1]=A_{1}\cup\dots\cup A_{N}$ (with $N$ small relative to sample size $n$), we use the available data estimate the error
$|\mathbb{E}[Y\mid f(X)\in A_{i}]-\mathbb{E}[f(X)\mid f(X)\in A_{i}]|,$
and aggregate over bins $i=1,\dots,N$, to estimate $\Delta_{\operatorname{ECE}}(f)$ (Błasiok et al., 2023). In fact, this type of binned approximation (known as the binned ECE) is actually estimating the ECE of the binned approximation to the function $f$ (that is, a function $g$ that is piecewise constant, taking the value $g(x)=\mathbb{E}[f(X)\mid f(X)\in A_{i}]$ for each $x\in A_{i}$)—and without further assumptions, it is possible to have $\Delta_{\operatorname{ECE}}(g)\approx 0$ even when $\Delta_{\operatorname{ECE}}(f)$ is large. This challenge is closely connected to our hardness result, Theorem 4.1.
@@ 5.1 Isotonic regression enjoys uniform asymptotic control of Cutoff Calibration Error
Suppose $(X_{i},Y_{i})\overset{iid}{\sim}P$ for $i=1,\dots,n$. Assume $Y_{i}\in[0,1]$ and $f$ is fixed. Let $h_{n}$ refer to running isotonic regression on $\{(f(X_{i}),Y_{i})\}_{i=1}^{n}$. Fix $\delta>0$. Then,
$\Delta_{\operatorname{Cutoff}}^{P}(h_{n}\circ f)\leq\frac{30+2\sqrt{2\log(2/\delta)}}{\sqrt{n}}$
@@ 5.2 Platt scaling does not necessarily calibrate, even in dCE
Platt scaling (Platt et al., 1999) is a popular post-hoc calibration method which essentially consists of running a univariate logistic regression of observed $Y$ values on previous forecasts f(X)f(X). [^6] For finite samples, Platt scaling minimizes logistic loss with $\tilde{Y}_{i}=Y_{i}\cdot\frac{\sum_{i=1}^{n}Y_{i}+1}{\sum_{i=1}^{n}Y_{i}+2}+(1-Y_{i})\frac{1}{\sum_{i=1}^{n}(1-Y_{i})+2}$ in place of $Y_{i}$ to improve empirical performance. The issue, in short, is that there is no guarantee that Platt scaling will converge asymptotically to a calibrated function. Since $\Delta_{\operatorname{dCE}}(f)=0$ if and only $f$ is perfectly calibrated, this implies that Platt scaling cannot provide strong asymptotic calibration on dCE, ECE, and Cutoff Calibration Error. In Figure 4, we exhibit an example of a data-generation process for which the population Platt scaling algorithm yields an $h:[0,1]\rightarrow[0,1]$ (attained by optimizing logistic loss in expectation instead of over samples) that has non-zero dCE: $\Delta_{\operatorname{dCE}}(h\circ f)>0$.
@@ 6 Discussion
There are several promising research directions and intriguing connections that stem from our findings. First, while our work focuses on the i.i.d. setting, we think our findings may hold ramifications for the branch of calibration research that focuses on sequential prediction and could potentially inform the design of new online calibration algorithms. Second, there may be connections between our work and Sahoo et al. (2021), who also focus on “threshold” decision rules, although in a different setting than us, where forecasters provide a CDF estimate for a real-valued output. Third, the monotonic constraint used in our decision-theoretic guarantees may provide intriguing results in other contexts, such as that of Roth and Shi (2024), who examine the impact of the dimensionality of a prediction space on swap regret and calibration. We also note that our focus on interval regions of the prediction space is related to the 1-dimensional special case in Roth and Shi (2024), which builds on the work in Noarov et al. (2023) in showing that actionability is related to performance within convex subsets of the prediction space. Finally, while the decision-theoretic guarantees of Cutoff Calibration Error are broadest when $Y\in\{0,1\}$, we believe that there may be a more general guarantee related to Cutoff Calibration Error when $Y\in[0,1]$.
@@ 7 Acknowledgments
The authors gratefully acknowledge the National Science Foundation via grant DMS-2023109. J.S. and R.F.B. were partially supported by the Office of Naval Research via grant N00014-24-1-2544. J.S. was also partially supported by the Margot and Tom Pritzker Foundation. Z.R. was supported by NSF grant DMS-2413135. R.M.W. was partially supported by the NSF-Simons National Institute for Theory and Mathematics in Biology (NITMB) via grants NSF DMS-2235451 and Simons Foundation MP-TMPS-00005320.
@@ Appendix A Relationships among calibration measures: proofs and examples
In this section, we show that in the case where the prediction $f(X)$ is a discrete random variable, dCE and ECE are essentially equivalent.
@@
    $\mathbb{E}[|f(X)-g(X)|]\leq\Delta_{\operatorname{dCE}}(f)+\epsilon.$
@@
    Next, let $w$ be any function taking values in [1,1][-1,1]. We need to bound $\mathbb{E}[w(f(X))\cdot(Y-f(X))]$. First, define
@@
    $\mathbb{P}\{w(f(X))\neq\tilde{w}(g(X))\}\leq\frac{1}{b_{f}}\cdot\mathbb{E}{[|f(X)-g(X)|]}\leq\frac{1}{b_{f}}\left(\Delta_{\operatorname{dCE}}(f)+\epsilon\right).$
    Finally, we bound ECE. We have
@@
For any function $f$,
$\Delta_{\operatorname{Cutoff}}(f)\leq\Delta_{\operatorname{ECE}}(f).$
@@
    This follows immediately from Section 2.1, which demonstrates that Cutoff Calibration Error and ECE are both instances of wCE and that the complexity class of weight functions for ECE is a superset of the complexity class for Cutoff Calibration Error. ∎
@@
For any function $f$,
$\Delta_{\operatorname{dCE}}(f)^{2}/9\leq\Delta_{\operatorname{Cutoff}}(f).$
@@
    $\psi(z):=\sum_{j=1}^{N}j{{\mathbf{1}_{\left\{z\in A_{j}\right\}}}},$
@@
    $\mathbb{E}[Y\mid g(X)]=\mathbb{E}[\mathbb{E}[Y\mid\psi(f(X))]\mid g(X)]=\mathbb{E}[g(X)\mid g(X)]=g(X).$
@@
    $\psi(f(X))=j\implies f(X),V\in A_{j}\implies|f(X)-V|\leq\mathrm{diam}(A_{j})=1/N.$
@@
$\mathbb{E}|g(X)-f(X)|\leq\mathbb{E}|g(X)-V|+\mathbb{E}|V-f(X)|\leq\mathbb{E}|g(X)-V|+1/N.$
@@
$\mathbb{E}|g(X)-V|$ $=\mathbb{E}|\mathbb{E}[Y-f(X)\mid\psi(f(X))]|$
$=\sum_{j=1}^{N}\mathbb{P}\{f(X)\in A_{j}\}|\mathbb{E}[(Y-f(X))\mid f(X)\in A_{j}]|$
$=\sum_{j=1}^{N}|\mathbb{E}[(Y-f(X)){{\mathbf{1}_{\left\{f(X)\in A_{j}\right\}}}}]|$
$\leq N\Delta_{\operatorname{Cutoff}}(f).$
@@
$\Delta_{\operatorname{dCE}}(f)\leq N\Delta_{\operatorname{Cutoff}}(f)+1/N.$
@@
$\Delta_{\operatorname{dCE}}(f)\leq 2\sqrt{\Delta_{\operatorname{Cutoff}}(f)}+\Delta_{\operatorname{Cutoff}}(f)\leq 3\sqrt{\Delta_{\operatorname{Cutoff}}(f)}.$
@@
    The first inequality follows immediately from observing that indicator functions over intervals have a total variation of at most 2 and that, if $g\in\mathcal{B}_{M}$, then $g\in\mathcal{B}_{M}$.
    For the second inequality, by Lemma A.3, $\exists g_{1},g_{2}:[0,1]\rightarrow[0,\frac{M}{2}+1]$ non-decreasing such that $g=g_{1}-g_{2}$.
    Define $\mathcal{G}:=\{g:[0,1]\rightarrow\mathbb{[}0,\frac{M}{2}+1]\mid g\text{ non-decreasing}\}$.
    Define $\mathcal{G}^{\prime}=\{g:[0,1]\rightarrow[0,1]\mid g\text{ non-decreasing}\}$.
@@
$\sup_{g\in\mathcal{B}_{M}}\mathbb{E}[(Y-f(X))g(f(X))]$
$\leq\sup_{g_{1},g_{2}\in\mathcal{G}}\left(\mathbb{E}[(Y-f(X))g_{1}(f(X))]-\mathbb{E}[(Y-f(X))g_{2}(f(X))]\right)$
$=\sup_{g\in\mathcal{G}}\mathbb{E}[(Y-f(X))g(f(X))]+\sup_{g\in\mathcal{G}}\mathbb{E}[(Y-f(X))(-g(f(X)))]$
$\leq 2\sup_{g\in\mathcal{G}\cup(-\mathcal{G})}\mathbb{E}[(Y-f(X))g(f(X))]$
$=2\sup_{g\in\mathcal{G}^{\prime}\cup(-\mathcal{G}^{\prime})}\mathbb{E}\left[(Y-f(X))(M/2+1)\cdot g(f(X))\right]$
$=(M+2)\sup_{g\in\mathcal{G}^{\prime}\cup(-\mathcal{G}^{\prime})}\mathbb{E}[(Y-f(X))g(f(X))]|$
$\leq(M+2)\Delta_{\operatorname{Cutoff}}(f).$
@@
$\mathcal{G}^{\prime}=\overline{\mathrm{conv}}\{g(x)={{\mathbf{1}_{\left\{x\in[t,1]\right\}}}}\mid t\in[0,1]\}$
@@
$\sup_{g\in\mathcal{G}^{\prime}\cup(-\mathcal{G}^{\prime})}\mathbb{E}[(Y-f(X))g(f(X))]$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ =\max\left\{\sup_{g\in\mathcal{G}^{\prime}}\mathbb{E}[(Y-f(X))g(f(X))],\sup_{g\in\mathcal{G}^{\prime}}\mathbb{E}[(Y-f(X))(-g(f(X)))]\right\}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ =\max\left\{\sup_{t}\mathbb{E}[(Y-f(X)){{\mathbf{1}_{\left\{f(X)\in[t,1]\right\}}}}],\sup_{t}\mathbb{E}[(Y-f(X))(-{{\mathbf{1}_{\left\{f(X)\in[t,1]\right\}}}})]\right\}$
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leq\sup_{t}|\mathbb{E}[(Y-f(X)){{\mathbf{1}_{\left\{f(X)\in[t,1]\right\}}}}]|.$
@@
Suppose $f:[0,1]\rightarrow[-1,1]$ and has total variation at most $M$. Then, $\exists f_{1},f_{2}:[0,1]\rightarrow[0,\frac{M}{2}+1]$ non-decreasing such that $f=f_{1}-f_{2}$.
@@
    To show $f_{1}$ is non-decreasing, let $x\leq y$, where $x,y\in[0,1]$.